# Nearest Neighbor CCP-Based Molecular Sequence Analysis

### v1

@@ VI Conclusion
In this study, we addressed the challenges of analyzing molecular sequence data, which involves a large number of sequences and complex protein structures. We proposed an efficient and fast method called Nearest Neighbor Correlated Clustering and Projection (CCP-NN). The CCP-NN method is based on the original Correlated Clustering and Projection (CCP) technique, but it incorporates an NN search for computing the density map and correlations. Through a series of experimental evaluations, we compared the performance of CCP and CCP-NN in classifying molecular sequences. The results demonstrated that CCP-NN outperforms CCP in terms of classification accuracy while also reducing computational runtime. Future work includes further enhancements to the CCP-NN framework, such as incorporating additional information or integrating it with other machine-learning methods for comprehensive analysis of sequences.

### v2

@@ VI Conclusion
In this study, we addressed the challenges of analyzing molecular sequence data, which involves a large number of sequences and complex protein structures. We proposed an efficient and fast method called Nearest Neighbor Correlated Clustering and Projection (CCP-NN). The CCP-NN method is based on the original Correlated Clustering and Projection (CCP) technique, but it incorporates an NN search for computing the density map and correlations. Through a series of experimental evaluations, we compared the performance of CCP and CCP-NN in classifying molecular sequences. The results demonstrated that CCP-NN outperforms CCP in terms of classification accuracy while also reducing computational runtime. Future work includes further enhancements to the CCP-NN framework, such as incorporating additional information or integrating it with other machine-learning methods for comprehensive analysis of sequences. Exploring other methods to determine the optimal number of neighbors is also an interesting future extension, which could include methods like adaptive neighbor selection based on local density estimation, the Elbow method, and Stability analysis. Future research aiming at biological interpretability, such as combining k-NN with interpretable models and using post-hoc interpretability frameworks like SHAP or LIME, may be compelling.